In chemical industries the following chemical reaction is used to produce ammonia in
large scale

N2+3H2 2NH3

Based on mole concept, calculate the mass of nitrogen gas and hydrogen gas required in
kilogram to produce 1000 kg of ammonia by using the above chemical equation
1
Expert's answer
2020-05-19T08:42:21-0400

N2+3H2= 2NH3

1) m(N2)=1000 kg * 28 g / 34 g = 823.5 kg

2) m(H2)=1000 kg * 6 g / 34 g = 176.47 kg